OpenStudy (anonymous):

The constant a certain spring of negligible mass is given by K = 1600 N / m. a) What should be the distance from the compression of this spring so that it stores potential energy equal to 3.20 J? b) You placed vertically one end of the spring on the ground. Drop on the spring book 1.20 kg from a height of 0.80 m above the top end of the spring. Calculate the distance this spring's maximum compression.

OpenStudy (rock_mit182):

Spring's potential energy given by: \[PE = \frac{ 1 }{ 2 }kx ^{2}

OpenStudy (rock_mit182):

\[\[PE = \frac{ 1 }{ 2 }kx ^{2}]\]

OpenStudy (whpalmer4):

\[PE = \frac{1}2kx^2\] :-)